# SceneLLM: Implicit Language Reasoning in LLM for Dynamic Scene Graph Generation

### v1

@@ 1 Introduction
Compared to static image SGG, the intricate spatio-temporal motion interaction among objects and model inference bias caused by the long-tail distribution of the dataset, place the task of dynamic SGG in a more challenging position [6]. Even though the promising application prospects of dynamic SGG have spurred a multitude of related exertions and constructive works in terms of spatial-temporal modeling [7, 2, 3, 8, 9, 10, 6] and unbiased design[11, 12, 13], significantly advancing the field of scene understanding, dynamic SGG is still an arduous yet immaturely addressed task [3].
@@ 2.1 Scene Graph Generation (SGG)
The scene graph represents a typical semantic graph structure that captures the relationships between objects within a scene, facilitating understanding and representation of visual information in a structured form [1]. Since its initial proposal for image retrieval tasks in [28], scene graphs have garnered significant interest in computer vision, leading to numerous advancements aimed at enhancing performance in tasks like visual question answering (VQA), image and video captioning, retrieval, and generation [29, 30, 31]. Recently, scene graph generation (SGG) has become particularly impactful in robot planning tasks, where visual scene comprehension is crucial. For instance, [5] highlights SGG’s role in robotic action sequencing, while [4] addresses robot planning under partial observability by integrating local scene graphs for improved global task execution. Additionally, [32] introduces a goal-oriented Programming by Demonstration approach, employing scene graphs to treat the robot as a functional operator. Consequently, the quest for high-quality image and video scene graph generation has become a focal area of research.
@@ 2.2 Large Language Models (LLMs)
Recently, increasingly powerful large language models (LLMs) like GPT-4 [14] and LLaMA3 [15] have emerged, achieving remarkable success across various NLP tasks. These tasks include question answering [37], multilingual translation [38], and so on. Benefiting from pre-training on vast text corpora, LLMs have developed a rich repository of tacit knowledge [39], enabling them to perform well even outside traditional NLP tasks. In computer vision, for example, LLMs are making strides in scene understanding and visual representation [17, 40]. Furthermore, a detailed survey [22] highlights that LLMs display enhanced reasoning abilities as model size increases, encompassing areas such as semantic [41], visual [42], and mathematical reasoning [43]. A recent study [44] also explores LLMs’ potential in control engineering, a domain that demands complex integration of mathematical and engineering principles. Building on these developments, our work leverages the robust reasoning capabilities of LLMs to advance scene understanding tasks, tapping into their broad knowledge and adaptability.
@@ 3.2 Video-to-Language (V2L) Mapping Module
Optimal Transport (OT) Scheme. Since the frame-level token is independent and only contains spatial information of its corresponding frame. To understand complex triplets related to high temporal information in dynamic scenes, we have to further process the frame-level token sequence $\{t_{\tau}\}_{\tau=1}^{T}$ for temporal consistency. For example, the triplet <person-hold-cup> exists in a static frame (scene), while with temporal information (dynamic scenes), it could be <person-drinking from-cup>. To this end, we propose to apply an OT scheme [46, 47] to transport semantic information from different frame-level tokens to equip the generated implicit linguistic signal with temporal information as Fig.4 shown. Various units in the learned codebook $\mathbb{C}$ contain different semantic information, our first step is to find out which units should be gathered to make the units in the updated codebook $\mathbb{C^{+}}$ with the temporal semantic property. From the information theory perspective, gathering discrete units can be viewed as a process of entropy reduction because of the lower uncertain information in codebook $\mathbb{C^{+}}$. On the other hand, considering the semantic representation capability of the codebook $\mathbb{C^{+}}$, a reasonable codebook size is also important. In short, we shall update for an optimal codebook $\mathbb{C^{+}}$ with lower entropy and reasonable size.
@@
Next, to formulate the objective function that minimizing the entropy of $\mathbb{C}_{k}^{+}$, following [40, 49, 46], we first set the transport matrix $P\in\mathbb{R}^{\Delta s\times(k\cdot\Delta s)}$ representing the allocation of the learned units $u$ to the new unit $u^{+}$, where the $(j,i)$-th element is $P(u_{j}^{+},u_{i})$, the distance matrix $D\in\mathbb{R}^{\Delta s\times(k\cdot\Delta s)}$ representing the cost of transport, where the $(j,i)$-th element is $\log P\left(u_{i}\mid u_{j}^{+}\right)$, and the $H(P)$ representing the entropy of the probability distribution $P(u_{j}^{+},u_{i})$. Therefore, according to Eq.4, our objective function is:
$\underset{\boldsymbol{P}\in\mathbb{R}^{\Delta s\times(k\cdot\Delta s)}}{\arg\min}H(\boldsymbol{P})+\sum_{j}\sum_{i}\boldsymbol{P}(j,i)\boldsymbol{D}(j,i)$ (5)
@@
Input : learned codebook $\mathbb{C}$, initial step $k=0$, increment $\Delta s$, initial updated codebook $\mathbb{C}_{0}^{+}=\mathbb{C}$, candidate codebooks $L$ = [$\mathbb{C}_{0}^{+}$],
initial codebook size $S_{c}=0$, Flag$=$True
@@ 3      $k\leftarrow k+1$;
4       $S_{c}\leftarrow k\times\Delta s$;
@@ 3.4 Training and Inference
Training. Our framework is optimized in two stages: the VQ-VAE training and the SceneLLM training. In the first stage, we pre-train the VQ-VAE model to encode and quantize visual features of objects. To achieve this, similar to [45], VQ-VAE is optimized to reconstruct visual features via encoding, quantization, and decoding process. The VQ-VAE loss function consists of 3 sub-losses, namely reconstruction loss for encoder-decoder learning, embedding loss for codebook learning, and commitment loss for limiting arbitrary growth of codebook volume. The loss function is defined as follows:
@@
where $sg[\cdot]$ is the stop-gradient operator (i.e. identity function) with zero partial derivates.
@@ 4.2 Implementation Details
Following TPT[6] and OED[50], we train a Transformer-based detector for object detection which is initialized from a COCO-pre-trained model and fine-tuned on the AG dataset. We set the latent feature dimension $l$ of VQ-VAE to 512, the number of codebook size $m$ to 512, and the weighting factor $\lambda$ in 7 to 0.02. Following [40], the auto-regressive model $\mathcal{M}$ is implemented as a one Convolutional Gated Recurrent Layer with a kernel size of (1,1). During the first training stage (VQ-VAE training), we optimize VQ-VAE for 300,000 iterations using the AdamW optimizer with an initial learning rate of 3e-4. In Eq. (3), we employ a bottom-up hierarchical clustering algorithm, with average linkage to measure the distance between between clusters, to construct the spatial relation $C$ between object features $F_{d}$. In addition, $GCN$ consists of two graph convolutional layers. We use frozen LLaMA-13B [15] as our LLM. During the second training stage (SceneLLM training), we first optimize the MLP, GCN, and SGG predictor for 30,000 iterations with an initial learning rate of 1e-5. Then, for the LoRA process, we fine-tune the LLM using the AdamW optimizer with an initial learning rate of 1e-5 for 50,000 iterations. The weighting factor $\alpha$ in Eq. (8) is set to 0.5. We conduct our experiments on Nvidia A5000 GPU.
@@ 4.4 Ablation Studies
The results are shown in Table 3. It can be observed that our method (w/ LLaMA) achieves better performance than the two variants, which demonstrates the effectiveness of leveraging a powerful LLM. In addition, the worse performance of w/ T5 shows a less powerful LLM (T5) may hard-to-hand relationships in complex scenes.
@@
Impact of the Optimal Transport Scheme. In SceneLLM, we design an optimal transport-based scheme to generate a temporal-consistent language signal. To validate this strategy, we test a variant called w/o OT, where we remove the OT scheme and directly feed the frame-level tokens $\{t_{\tau}\}_{\tau=1}^{T}$ (with prompt texts) into the LLM. The declined performance of the w/o OT variant shown in Table 5 demonstrates the efficacy of our OT scheme, which reconstructs LLM-friendly language signal from frame-level tokens.

### v2

@@ 1 Introduction
Compared to static image SGG, the intricate spatio-temporal motion interaction among objects and model inference bias caused by the long-tail distribution of the dataset, place the task of dynamic SGG in a more challenging position. Even though the promising application prospects of dynamic SGG have spurred a multitude of related exertions and constructive works in terms of spatial-temporal modeling [6, 2, 3, 7, 8, 9] and unbiased design[10, 11, 12], significantly advancing the field of scene understanding, dynamic SGG is still an arduous yet immaturely addressed task [3].
@@ 2.1 Scene Graph Generation (SGG)
The scene graph represents a typical semantic graph structure that captures the relationships between objects within a scene, facilitating understanding and representation of visual information in a structured form [1]. Since its initial proposal for image retrieval tasks in [27], scene graphs have garnered significant interest in computer vision, leading to numerous advancements aimed at enhancing performance in tasks like visual question answering (VQA), image and video captioning, retrieval, and generation [28, 29, 30, 31, 32, 33]. Recently, scene graph generation (SGG) has become particularly impactful in robot planning tasks, where visual scene comprehension is crucial. For instance, [5] highlights SGG’s role in robotic action sequencing, while [4] addresses robot planning under partial observability by integrating local scene graphs for improved global task execution. Additionally, [34] introduces a goal-oriented Programming by Demonstration approach, employing scene graphs to treat the robot as a functional operator. Consequently, the quest for high-quality image and video scene graph generation has become a focal area of research.
@@ 2.2 Large Language Models (LLMs)
Recently, increasingly powerful large language models (LLMs) like GPT-4 [13] and LLaMA3 [14] have emerged, achieving remarkable success across various NLP tasks. These tasks include question answering [39], multilingual translation [40], and so on. Benefiting from pre-training on vast text corpora, LLMs have developed a rich repository of tacit knowledge [41], enabling them to perform well even outside traditional NLP tasks. In computer vision, for example, LLMs are making strides in scene understanding and visual representation [16, 42]. Furthermore, a detailed survey [21] highlights that LLMs display enhanced reasoning abilities as model size increases. A recent study [43] also explores LLMs’ potential in control engineering, a domain that demands complex integration of mathematical and engineering principles. Building on these developments, our work leverages the robust reasoning capabilities of LLMs to advance scene understanding tasks, tapping into their broad knowledge and adaptability.
@@ 3.2 Video-to-Language (V2L) Mapping Module
Optimal Transport (OT) Scheme. Since the frame-level token is independent and only contains spatial information of its corresponding frame. To understand complex triplets related to high temporal information in dynamic scenes, we have to further process the frame-level token sequence $\{t_{\tau}\}_{\tau=1}^{T}$ for temporal consistency. For example, the triplet <person-hold-cup> exists in a static frame (scene), while with temporal information (dynamic scenes), it could be <person-drinking from-cup>. To this end, we propose to apply an OT scheme [45, 46] to transport semantic information from different frame-level tokens to equip the generated implicit linguistic signal with temporal information as shown in Fig.4. Various units in the learned codebook $\mathbb{C}$ contain different semantic information, our first step is to find out which units should be gathered to make the units in the updated codebook $\mathbb{C^{+}}$ with the temporal semantic property. From the information theory perspective, gathering discrete units can be viewed as a process of entropy reduction because of the lower uncertain information in codebook $\mathbb{C^{+}}$. On the other hand, considering the semantic representation capability of the codebook $\mathbb{C^{+}}$, a reasonable codebook size is also important. In short, we shall update for an optimal codebook $\mathbb{C^{+}}$ with lower entropy and reasonable size.
@@
$=-\sum_{u_{j}^{+}\in\mathbb{C}_{k}^{+}}\sum_{u_{i}\in\mathbb{C}}P\left(u_{j}^{+},u_{i}\right)\log P\left(u_{j}^{+}\right)$
$\quad-\sum_{u_{j}^{+}\in\mathbb{C}_{k}^{+}}\sum_{u_{i}\in\mathbb{C}}P\left(u_{j}^{+},u_{i}\right)\left(-\log\frac{P(u_{j}^{+},u_{i})}{P(u_{j}^{+})}\right)$
@@
Next, to formulate the objective function that minimizing the entropy of $\mathbb{C}_{k}^{+}$, following [42, 48, 45], we first set the transport matrix $P\in\mathbb{R}^{m\times(k\cdot\Delta s)}$ representing the allocation of the learned units $u$ to the new unit $u^{+}$, where the $(j,i)$-th element is $P(u_{j}^{+},u_{i})$, the distance matrix $D\in\mathbb{R}^{m\times(k\cdot\Delta s)}$ representing the cost of transport, where the $(j,i)$-th element is $\log P\left(u_{i}\mid u_{j}^{+}\right)$, and the $H(P)$ representing the entropy of the probability distribution $P(u_{j}^{+},u_{i})$. Therefore, according to Eq.4, our objective function is:
$\underset{\boldsymbol{P}\in\mathbb{R}^{m\times(k\cdot\Delta s)}}{\arg\min}H(\boldsymbol{P})+\sum_{j}\sum_{i}\boldsymbol{P}(j,i)\boldsymbol{D}(j,i)$ (5)
@@
Input : learned codebook $\mathbb{C}$, initial step $k=0$, increment $\Delta s$, initial updated codebook $\mathbb{C}_{0}^{+}=\mathbb{C}$, candidate codebooks $L$ = [$\mathbb{C}_{0}^{+}$], Flag$=$True
@@ 3      $k\leftarrow k+1$;
4       Expand codebook size to $k\times\Delta s$;
@@ 3.4 Training and Inference
Training. Our framework is optimized in two stages: the VQ-VAE training and the SceneLLM training. At the first stage, we pre-train the VQ-VAE model to encode and quantize visual features of objects. To achieve this, similar to [44], VQ-VAE is optimized to reconstruct visual features via encoding, quantization, and decoding process. The VQ-VAE loss function consists of 3 sub-losses, namely reconstruction loss for encoder-decoder learning, embedding loss for codebook learning, and commitment loss for limiting arbitrary growth of codebook volume. The loss function is defined as follows:
@@
where $sg[\cdot]$ is the stop-gradient operator (i.e. identity function) with zero partial derivatives.
@@ 4.2 Implementation Details
Following OED[49], we train a Transformer-based detector for object detection which is initialized from a COCO-pre-trained model and fine-tuned on the AG dataset. We set the latent feature dimension $l$ of VQ-VAE to 512, the number of codebook size $m$ to 512, and the weighting factor $\lambda$ in 7 to 0.02. Following [42], the auto-regressive model $\mathcal{M}$ is implemented as a one Convolutional Gated Recurrent Layer with a kernel size of (1,1). During the first training stage (VQ-VAE training), we optimize VQ-VAE for 300,000 iterations using the AdamW optimizer with an initial learning rate of 3e-4. In Eq. (3), we employ a bottom-up hierarchical clustering algorithm, with average linkage to measure the distance between between clusters, to construct the spatial relation $C$ between object features $F_{d}$. In addition, $GCN$ consists of two graph convolutional layers. We use frozen LLaMA-13B [14] as our LLM. During the second training stage (SceneLLM training), we first optimize the MLP, GCN, and SGG predictor for 30,000 iterations with an initial learning rate of 1e-5. Then, for the LoRA process, we fine-tune the LLM using the AdamW optimizer with an initial learning rate of 1e-5 for 50,000 iterations. The weighting factor $\alpha$ in Eq. (8) is set to 0.5. We conduct our experiments on Nvidia A5000 GPU.
@@ 4.3 Performance Results and Comparison
4.4 Visualization
Fig 5: Visualization results of our method. All results are given under the SGDET setup. (Zoom in for the best view.)
As shown in Fig.5, SceneLLM’s visual detection results in dynamic scenes highlight its superior ability to model spatiotemporal relationships. This advancement is the result of SceneLLM’s strong inference capabilities on fine-grained spatiotemporal dependencies and the vast and rich pre-trained knowledge contained in LLMs, which enables the ability to accurately parse complex real-world scenes, even as object-to-human interactions dynamically change over time.
@@ 4.5 Ablation Studies
Fig 6: Impact of weight factor $\alpha$ (With Constraint)
Fig 7: Impact of weight factor $\alpha$ (Without Constraint)
@@
Impact of weight factor $\alpha$. We obtained the optimal weight factor $\alpha$ through grid search. As shown in Fig.6 and Fig.7, under the R@K (K=[10,20,50]) metrics, the model achieved the best performance when alpha was set to 0.5, regardless of constrained or unconstrained settings.
@@
The results are shown in Table 3. It can be observed that our method (w/ LLaMA) achieves better performance than the two variants, which demonstrates the effectiveness of leveraging a powerful LLM. In addition, the worse performance of w/ T5 shows that a less powerful LLM (T5) may hard to handle relationships in complex scenes.
@@
Impact of the Optimal Transport Scheme. In SceneLLM, we design an optimal transport-based scheme to generate a temporally consistent language signal. To validate the effectiveness of our strategy, we test three variants: 1) w/o OT, where we remove the Optimal Transport (OT) scheme and directly feed the frame-level tokens ${t_{\tau}}_{\tau=1}^{T}$ (along with the prompt texts) into the LLM; 2) Temporal Convolution (TC), where we replace the OT scheme with a temporal convolution module to capture temporal correlations between frame-level tokens before feeding them into the LLM; 3) Clustering, where we substitute the OT scheme with a clustering algorithm to aggregate tokens with temporal correlations and then derive a new codebook $\mathbb{C}^{+}$. As shown in Table 5, our method outperforms these variants, demonstrating the efficacy of the OT scheme in reconstructing an LLM-friendly language signal from frame-level tokens.